OpenStudy (anonymous):

if i have a parallelogram that has 2 diagonals one side is 4x+5 the other side is 28 what is the value of x?

OpenStudy (anonymous):

x= 23/4

OpenStudy (anonymous):

what did you do to get that

jhonyy9 (jhonyy9):

4x+5=28 because every two diagonals are the same longs hence from this x=23/4